When a liquid is introduced between the length and the plate in the NEWTON'S ring Apparatus the diameter of the 15th ring changes from 2cm to 1.8 CM. The refractive index of a liquid is
1
Expert's answer
2020-09-01T10:37:05-0400

Solution

Given data in the question

Order of newton ring (n) =15

Before introducing any liquid between

the length and the plate in the Newton's ring Apparatus there was air so it's Diameter "(D_n) _{air}=2 cm"

"=2\times10^{-2}m"

After introduced liquid diameter

"(D_n) _{liquid}=1.8cm"

"=1.8\times10^{-2}m"

Refractive index of this liquid can be written as

"\mu=\frac{(D_n) ^2_{air}}{(D_n)^2 _{liquid}}"


"\mu=\frac{(2\times10^{-2}) ^2}{(1.8\times10^{-2}) ^2}"

"\mu=\frac{4}{3.24}"


"\fcolorbox{red}{yellow}{$\mu=1.23$}"

Therefore refractive index of liquid is 1.23 .